OpenStudy (anonymous):

If two brothers work together they can change the oil in the family car in three minutes. If it takes the older brother 12 minutes to change the oil alone, how long does it take for the younger brother to change the oil?

OpenStudy (raden):

(O * Y)/(O + Y) = 3 (12 * Y)/(12 + Y) = 3 solve for Y

OpenStudy (raden):

Y = Young O = old

OpenStudy (anonymous):

In these kind of sums always find the work done in unit time in 1 min both can do only 1/3 part of work but in 1 min older does alone 1/12 part of work thus in 1min younger does 1/3-1/12=(4-1)/12=3/12 =1/4 of work thus it takes 4mins for the younger to change the oil